OpenStudy (anonymous):

2x^3/6x^2-9, is this even, odd or neither?

OpenStudy (campbell_st):

do you know the rules for odd and even functions...

OpenStudy (campbell_st):

and it doesn't depend on the value of x.

OpenStudy (anonymous):

It just asks is this even, odd or neither

OpenStudy (anonymous):

Even and odd functions

OpenStudy (campbell_st):

ok... an even function occurs when f(-x) = f(x) and odd function occurs when f(-x) = -f(x) if neither condition is met, the the function is neither odd nor even. So substitute x = -x into the function and see what you get.. post the answer

OpenStudy (anonymous):

First step. 2x^3/6x^2-9 -6x^2 +9

OpenStudy (campbell_st):

close, it's \[f(-x) = \frac{2(-x)^3}{6(-x)^2 - 9}\] can you simplify this..?

OpenStudy (anonymous):

(X-3)^2

OpenStudy (campbell_st):

not quite \[(-x)^3 = -x^3\] and \[(-x)^2 = x^2\] does that make sense..?

OpenStudy (campbell_st):

ok.... so then simplifying you can say \[f(-x) = \frac{-2x^3}{6x^2 - 9}\] which is \[f(-x) = -\left( \frac{2x^3}{6x^2 - 9} \right)\] does that look familiar.... and please scroll up and check to conditions for odd, even and neither functions.

OpenStudy (anonymous):

I am confused on how to solve this.

OpenStudy (anonymous):

Would it be -2^3/-6x^2+9

OpenStudy (campbell_st):

you don't solve it the process is to 1. replace x with -x 2. simplify where necessary 3. compare it to the original function so see if it's odd or even or neither

OpenStudy (anonymous):

It would be neither

OpenStudy (campbell_st):

ok... so when I subsituted x = -x, and simplified I got \[f(-x) = - \left( \frac{2x^3}{6x^2 - 9} \right)\] which means \[f(-x) = - f(x)\] which is the condition for an odd function

OpenStudy (anonymous):

F(-x)= 2(-x)^3 turns into -2x^3 then 6(-x)^2 -9 tuns into 6x^2 -9 ; the top number is not the same. For -f(x) = -(2^3/6x^2-9) = -2x^3/-6x^2+9 so it would be odd rather than neither. Thanks!

OpenStudy (campbell_st):

that is incorrect... if you \[-1 \times \frac{2x^3}{6x^2 - 9} = \frac{-2x^3}{6x^2 - 9}\] the denominator is not multiplied by -1, just the numerator... its like saying what is \[-1 \times \frac{1}{2} = \frac{-1}{2}... or \frac{1}{-2}\] but remembering the rules for multiplying and dividing negatives \[\frac{-1}{-2} = \frac{1}{2}\] which can't occur if you multiply 1/2 by -1

OpenStudy (anonymous):

Got it!